20220515_DX_Day. DX ASM

20:23 (UT) testing magnets

20:31 (UT) nominal flat applied with 554 modes limiting actuator 440 (in the first ring from the center), but the force there is below 0.3N RB temp 13.8

20:35 (UT) re-applied Default flat /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav and we 568 modes, limiting actuator 440

20:37 (UT) testing magnet of actuator 581

  • -0.045 62.81
  • +10nm
  • -0.053 62.82
  • +10nm
  • -0.063 62.83
  • +10 nm
  • -0.070 62.84
  • +10 nm
  • -0.079 62.84
  • +10 nm
  • -0.089 62.87
  • -10 nm
  • -0.079 62.84
  • -10 nm
  • -0.079 62.84
  • -10 nm
  • -0.064 62.82
  • -10 nm
  • -0.054 62.81
  • -10 nm
  • -0.045 62.81
  • -10 nm
  • -0.038 62.80
  • -10 nm
  • -0.032 62.78
  • -10 nm
  • -0.023 62.78
  • -10 nm
  • -0.023 62.76
  • -10 nm
  • -0.009 62.75
  • -10 nm
  • +0.002 62.76
  • -10 nm
  • +0.010 62.74
  • -10 nm
  • +0.01 62.72

20:59 (UT) the magnet of 581 seems to be OK

20:59 (UT) resting and setting again Default flat /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav and we 554 modes, limiting actuator 440

21:00 (UT) testing magnet of actuator 580

  • 63.78 -0.075
  • +20 nm
  • 63.80 -0.09
  • +20 nm
  • 63.82 -0.11
  • -20 nm
  • 63.80 -0.09
  • -20 nm
  • 63.79 -0.075
  • -20 nm
  • 63.76 -0.06
  • -20 nm
  • 63.75 -0.045
  • -20 nm
  • 63.72 -0.03
  • -20 nm
  • 63.70 -0.01
  • - 20 nm
  • 63.68 -0.00
  • -20 nm
  • 63.66 +0.02

21:14 (UT) the magnet of 580 seems to be OK

20:15 (UT) resting and setting again Default flat /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav and we 554 modes, limiting actuator 440

21:00 (UT) testing magnet of actuator 582

  • 62.73 -0.00
  • +20 nm
  • 62.75 -0.02
  • +20 nm
  • 62.78 -0.03
  • +20nm
  • 62.80 -0.05
  • +20 nm
  • 62.81 -0.06
  • +20 nm
  • 62.83 -0.07
  • -20 nm
  • 62.82 -0.06
  • -20 nm
  • 62.80 -0.05
  • -20 nm
  • 62.77 -0.03
  • -20 nm
  • 62.74 -0.02
  • -20 nm
  • 62.74 -0.00
  • -20 nm
  • 62.71 +0.01
  • -20 nm
  • 62.69 +0.02

21:25 (UT) the magnet of 582 seems to be OK

21:30 (UT) resting and setting again Default flat /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav and we 554 modes, limiting actuator 217 RB temp 14.2

21:31 (UT) testing magnet of actuator 579 (note 599 next shows some jumps)

  • 64.387 +0.064
  • -20 nm
  • 64.376 +0.079
  • -20 nm
  • 64.353 +0.095
  • +20 nm
  • 64.373 +0.082
  • +20 nm
  • 64.394 +0.064
  • + 20 nm
  • 64.421 +0.048
  • +20 nm
  • 64.431 +0.034
  • +20 nm
  • 64.465 +0.015
  • +20 nm
  • 64.472 -0.000
  • +20 nm
  • 64.501 -0.016

21:42 (UT) the magnet of 579 seems to be OK

21:42 (UT) resting and setting again Default flat /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav and we 554 modes, limiting actuator 217 RB temp 14.3

21:44 (UT) testing magnet of actuator 583

  • 65.433 -0.056
  • +20 nm
  • 65.434 -0.068
  • +20 nm
  • 65.472 -0.079
  • +20 nm
  • 65.497 -0.089
  • +20 nm
  • 65.512 -0.099
  • -20 nm
  • 65.489 -0.090
  • -20 nm
  • 65.477 -0.080
  • -20 nm
  • 65.454 -0.070
  • -20 nm
  • 65.435 -0.060
  • -20 nm
  • 65.396 -0.050
  • -20 nm
  • 65.389 -0.039
  • -20 nm
  • 65.361 -0.032
  • -20 nm
  • 65.354 -0.022
  • -20 nm
  • 65.327 -0.010
  • -20 nm
  • 65.314 -0.001
  • -20 nm
  • 65.281 +0.008
  • -20 nm
  • 65.262 +0.018

22:00 (UT) the magnet of 583 seems to be OK

22:01 (UT) resting and setting again Default flat /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav and we 554 modes, limiting actuator 217 RB temp 14.4

22:02 (UT) shell rested

22:41 (UT) starting AO tests

23:01 (UT) setting Default flat /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav and we 554 modes, limiting actuator 217 RB temp 14.6, actuators 357, 244, 468, 599 are the 'classic' checkerboard pattern, we should consider putting them in the MMC list.

23:17 (UT) Problems acquiring the ARGOS source

23:23 (UT) bin 1, 900 Hz, 405 modes, no disturbance, AO loop gain zero

23:25 (UT) 232554, 233607, 232619

23:39 (UT) all three data sets show normal noise (~5nm rms) from the ASM position sensors no peaks around 130-140 Hz

23:57 (UT) still trying to load disturbance

00:17 (UT next day) we were finally able to load the disturbance

00:18 (UT) bin 1, 900 Hz, 405 modes, with disturbance, AO loop gain zero

00:19 (UT) 001554, 001627, 001640 with 1.6" seeing (0.8 disturbance) and zero gain

00:21 (UT) 001852, 001913, 001929 with some minimal gains and disturbance

00:28 (UT) waiting for OCAM to cool down

00:41 (UT) running AO loop optimization (we have some coma that is not offloaded)

00:46 (UT) completed optimization

00:46 (UT) Optimized with 0.8 Disturbance (1.6” seeing ARGOS) 004607, 004642, 004655

-- GuidoBrusa - 15 May 2022
Topic revision: r4 - 16 May 2022, GuidoBrusa
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback